The refractive index of a transparent material can be determined by measuring the critical angle when the solid is in air. If θc= 41.1° what is the index of refraction of the material? Incorrect. Tries 1/99 Previous Tries

A light ray strikes this material (from air) at an angle of 33.0° with respect to the normal of the surface. Calculate the angle of the reflected ray (in degrees). Tries 0/99

Calculate the angle of the refracted ray (in degrees). Tries 0/99

Assume now that the light ray exits the material. It strikes the material-air boundary at an angle of 33.0° with respect to the normal. What is the angle of the refracted ray?
Answer #1

Using Snell's law,

n2/n1 = sin(theta_1)/sin(theta_2)

A)

n/1 = sin(90)/sin(41.1)

n = 1.52

B)

Angle of reflection = 33°

C)

1.52/1 = sin(33)/sin(theta)

Theta = 21°

D)

1.52/1 = sin(theta)/sin(33)

Theta = 55.88°

Comment in case any doubt please rate my answer....